# Probing Deep into Temporal Profile Makes the Infrared Small Target Detector Much Better

### v1

@@ Abstract
Infrared small target (IRST) detection is challenging in simultaneously achieving precise, universal, robust and efficient performance due to extremely dim targets and strong interference. Current learning-based methods attempt to leverage “more” information from both the spatial and the short-term temporal domains, but suffer from unreliable performance under complex conditions while incurring computational redundancy. In this paper, we explore the “more essential” information from a more crucial domain for the detection. Through theoretical analysis, we reveal that the global temporal saliency and correlation information in the temporal profile demonstrate significant superiority in distinguishing target signals from other signals. To investigate whether such superiority is preferentially leveraged by well-trained networks, we built the first prediction attribution tool in this field and verified the importance of the temporal profile information. Inspired by the above conclusions, we remodel the IRST detection task as a one-dimensional signal anomaly detection task, and propose an efficient deep temporal probe network (DeepPro) that only performs calculations in the time dimension for IRST detection. We conducted extensive experiments to fully validate the effectiveness of our method. The experimental results are exciting, as our DeepPro outperforms existing state-of-the-art IRST detection methods on widely-used benchmarks with extremely high efficiency, and achieves a significant improvement on dim targets and in complex scenarios. We provide a new modeling domain, a new insight, a new method, and a new performance, which can promote the development of IRST detection. Codes are available at https://github.com/TinaLRJ/DeepPro.
@@ 1 Introduction
The objective of Infrared Small Target (IRST) detection [1, 2] is to identify and localize moving targets that are typically “small” and “dim” 111“Small” means the target usually occupies a small portion of the overall image, and “dim” indicates that the target exhibits a weak signal intensity and low contrast relative to its background. from infrared images. It has been a longstanding, fundamental yet challenging problem in infrared remote sensing [3, 4, 5], and has broad applicability in both civilian and military domains, including reconnaissance, precision guidance [6, 7], maritime surveillance [8, 9], early fire warnings [10, 11], vehicle detection [12, 13], and UAV monitoring [14, 15]. In recent years, IRST detection has attracted increasing attention and made measurable progress over the past few decades; however, most IRST detection methods have not been capable of simultaneously achieving outstanding performance in precision, robustness, efficiency, and universality due to the following challenges [11, 16, 17, 18].
  • High-precision related challenges. On the one hand, the targets themselves are usually very small (e.g., typically occupy $\leq 9\times 9$ pixels), dim (e.g., with an SNR $\leq 3$) [7, 19, 20], and with insufficient appearance features, such as shape and texture information. Such characteristics increase the difficulty of developing effective feature representations. On the other hand, the background is usually nonsmooth and nonuniform, containing clutter and noise. Clutter refers to imaged real things, while noise refers to electronic noise and inaccuracies of the signal processor. They almost occupy the entire image, and some of them are strong and dynamic, similar to the targets. This complicates the effective suppression of false alarms.
  • High-robustness related challenges. Apart from the aforementioned challenges, the target characteristics evolve temporally under motion conditions. The intensity of the target varies, and the local background neighborhood of the target also changes dynamically, making it difficult to detect the target continuously.
@@
  • High-university related challenges. In practical applications, there are various targets, such as air-breathing aircraft, rocket-propelled vehicles, and unpowered aerodynamic gliders [21, 22], each exhibiting distinct infrared radiation characteristics. In addition, many complex backgrounds (e.g., sea, cloud, and land environments) present unique clutter properties. Adapting to different targets and backgrounds is challenging.
@@
Therefore, in diverse complex scenarios, which modeling domain can better exhibit the targets? Which prior information can more effectively capture the essential differences between target and background? Exploring these questions is critical to achieve precise, robust, efficient, and universal detection.
@@
Existing deep learning-based methods focus on extracting “more” information to improve the detection performance. For example, many single-frame methods [23, 19, 24] extensively explore the local saliency and global information in the spatial domain via the complex network structure. Besides, some multi-frame methods [25, 26, 27, 28] explore motion information in the short-term spatial-temporal domain. Although performance improvements are achieved in some scenarios, these methods suffer from severe performance degradation under complex conditions. This comes from the limitation that the target information is weak and that other discriminative information is insufficient in both domains, as illustrated in Fig. 1(a) and (b). In addition, extracting “more” information inevitably expands receptive fields, causing severe computational redundancy that fundamentally undermines real-time processing capabilities. Therefore, given our previous analysis, the focus of this work is to explore the “more essential” information from a more crucial profile and develop a highly effective and efficient method for IRST detection, especially under complex conditions.
@@
Figure 1: The typical difficulties in IRST detection by observing targets from different domains. (a) Spatial domain: dim and small targets may be nearly unobservable; (b) Short-term spatial-temporal domain: the targets are indistinguishable from the interference similar to the targets; (c) Temporal profile: this profile records the dynamic statistical details of all signals on the same spatial location. Dim and small target signals are complete and more prominent in the temporal profile. The temporal profile contains more essential information.
@@
In this paper, we study the inherent characteristics of IRST detection task in depth. We reveal that the temporal profile is crucial especially under complex conditions, which dramatically amplifies target saliency as shown in Fig. 1(c). Through further theoretical analysis, we find that the temporal profile contains global temporal saliency information of target signals and correlation information between different signals, which are more essential for IRST detection. To validate the importance of the temporal profile in well-trained networks, we built the first prediction attribution [29, 30] tool that can analyze which input pixels impact the predictions of IRST detection networks. The attribution results show that the most influential pixels are concentrated in a small area centered on the target and continuously distributed along the time axis, which illustrates the importance of the temporal profile.
@@
Inspired by the above research, we remodel the IRST detection task as a one-dimensional signal anomaly detection task in the temporal profile, where targets manifest as statistical outliers in time-sequential signals. Then, we designed a pixel-wise temporal probe mechanism (TPro) to extract essential information in the temporal profile. Based on TPro, we propose an effective and efficient deep temporal probe network (i.e., DeepPro) for IRST detection, which performs multiplications and additions only in the time dimension. We conducted numerous experiments on widely-used benchmarks. As shown in Fig. 2, our DeepPro achieves significantly superior performance as compared to existing state-of-the-art methods employing complex spatial-temporal computations. Meanwhile, our DeepPro exhibits higher robustness on dim targets and in complex scenarios with strong noise, achieving a 78.4% reduction in parameter count and over 2×2$\times$2 × acceleration improvement as compared to the lightest and fastest single-frame IRST (SIRST) detection network (i.e., Res-UNet [31]).
@@
  • Inspired by our research, we remodel the IRST detection task as a one-dimensional signal anomaly detection task. Then, we propose an IRST detection network (i.e., DeepPro) to leverage the essential temporal profile information with multiply-add operations only in the time dimension.
  • Experimental results show that DeepPro not only achieves a significant performance improvement in IRST detection with extremely high efficiency and continuity, but also has high robustness to dim targets and scenes with strong noise.
@@ 2.1 Single-frame infrared small target detection
Traditional SIRST detection methods mainly include spatial-filtering-based methods [38, 39, 40, 41, 42], sparsity-and-low-rank-based methods [43, 44, 45, 46, 47, 48], and human-visual-system-based (HVS) methods [49, 50, 51, 52, 53, 54]. Recently, deep learning-based methods achieved better detection performance than traditional ones due to their strong feature learning abilities. Drawing inspiration from traditional methods, some networks are integrated with some prior knowledge for robust SIRST detection, such as the local spatial saliency differences between targets and backgrounds [55, 56, 57], the sparsity of targets and the low rankness of backgrounds [58, 37, 59]. Dai et al. [55] incorporated local contrast measure [52] into the network. Zhang et al. [56] converted the local contrast measure into a nonlocal quadrature difference measure in deep feature space. Zhou et al. [58] developed a deep unrolling network to exploit the low-rankness of backgrounds and the sparsity of targets. Liu et at. [59] combined low rankness and local smoothness prior of IRSTs to better model the background. Wu et al. [37] formulated the IRST detection task as sparse target extraction, low-rank background estimation, and image reconstruction in a relaxed robust principle component analysis model. For performance improvements, some methods combine IRST properties with the attention mechanism [60] to learn correlations among pixels and pay attention to key features. Li et al. [34] cascaded channel and spatial attention module to enhance multi-level features. Zhang et al. [61] proposed a Taylor finite difference-inspired edge block and a two-orientation attention aggregation block to highlight the shape of the targets. Liu et al. [19] adopted the self-attention mechanism of Transformer to learn the non-local correlation of image features. Beside feature extraction, Liu et al. [59] explored the loss functions of IRST detection and proposed a scale and location sensitive (SLS) loss to boost detection performance. Moreover, some researchers investigated the weakly-supervised detection [62, 63, 64, 65], in which the pseudo-masks are generated based on the local saliency of the IRST. SIRST detection is applied to locate the salient targets, but is not suitable for targets with low SNR.
@@ 2.2 Multi-frame infrared small target detection
MIRST detection aims at exploiting both spatial and temporal information from input image tensors, and is expected to achieve better and more robust performance than SIRST detection. Existing MIRST detection methods can be categorized into filtering-based methods [66, 67, 68, 69, 70, 71], sparsity-and-low-rank-based methods [72, 73, 74, 33, 8, 75, 76], and deep learning-based methods [77, 26, 7, 27, 78, 79, 80].
Filtering-based methods are the earlier developed methods, which mainly uses the temporal information among frames. Reed et al. [66, 67, 68] proposed a series of 3D matched filtering methods to highlight the targets through spatial filtering and temporal energy accumulation. Gao et al. [70] proposed a time variance and patch contrast filter to improve the local contrast of targets via multi-scale patch-based contrast measure. Li et al. [71] achieved MIRST detection by using the directional morphological filtering and spatial-temporal semantic information (e.g., the consistency of targets and the fluctuation of clutters).
@@ 3.1.1 Global temporal saliency
We analyzed the waveform characteristics of target signals in the temporal profile. Given that the velocity of small targets exhibits minimal variation over the short time they pass through a detection unit completely, we modeled the moving targets in uniform linear motions. In Fig. 3, we display some toy examples of the temporal profiles of the targets with different intensities, sizes, and speeds. The fluctuation of a target signal $G$ is regular with a single up-down process. As a target moves into and out of a detection unit, the target intensity in this unit first increases and then decreases. The amplitude of $G$ is related to the maximum intensity of the part of the target passing through the detection unit, and a brighter target usually has a higher peak value. Besides, the width of $G$ is related to the size and speed of the target. Bigger targets typically exhibit wider signals on the temporal profile, while targets with higher speeds under the same conditions produce narrower signals. Therefore, assuming that the fluctuations of other signals can be ignored, the target signals are saliency in the temporal profile.
Figure 4: Spatial and temporal-profile visualizations of targets in real scenes with noise and clutter. The red annotation marks the target.
@@ 3.1.2 Correlation information
As the intensities of noise and clutter signals increase, the global temporal saliency becomes weak and damaged (e.g., Fig. 4(d)). Therefore, we deeply extract other essential distinctive characteristics between noise $N$, clutter $B$, and target $G$ in the temporal profile.
@@
[…]
@@
Then, to investigate the differences between target and clutter, we mixed target and different clutters (e.g., land clutter and cloud clutter), and analyzed the correlation function of the mixed signals (i.e., $f_{G+B}(t)$). At first, we got different clutters from the real scenarios. After decentralizing and scaling, we added the fixed-intensity target signal to a series of clutter signals with different standard deviations $\sigma_{B}$ [^3] Note that the standard deviation of clutter in the time dimension is different from its fluctuation intensity in the space dimension. Usually, the temporal standard deviation of a clutter is smaller than its local spatial fluctuation since the background is slowly changing., as shown in Fig. 5(b1) and (c1). The typical clutters usually include low-frequency signals (e.g., slowly-changing cloud in c1) and high-frequency signals (e.g., camera blind elements and flickering background in b1). High-frequency signals are more likely to be falsely detected as targets due to their spatial and temporal saliency. When the clutter signal in the time dimension is not strong, there are some obviously distinctive characteristics between targets and clutters. Usually, the fluctuations of the clutter are dense, and they are correlated with each other on the amplitude in a period but uncorrelated on the occurrence moment. However, these characteristics are still inadequate for the detection when the clutter has a high standard deviation. Following the above analysis, we calculated the correlation function $R_{G,B}(\tau)$ of $f_{G+B}(t)$, i.e.,
[…]

### v2

@@ Abstract
Infrared small target (IRST) detection is challenging in simultaneously achieving precise, robust, and efficient performance due to extremely dim targets and strong interference. Current learning-based methods attempt to leverage “more” information from both the spatial and the short-term temporal domains, but suffer from unreliable performance under complex conditions while incurring computational redundancy. In this paper, we explore the “more essential” information from a more crucial domain for the detection. Through theoretical analysis, we reveal that the global temporal saliency and correlation information in the temporal profile demonstrate significant superiority in distinguishing target signals from other signals. To investigate whether such superiority is preferentially leveraged by well-trained networks, we built the first prediction attribution tool in this field and verified the importance of the temporal profile information. Inspired by the above conclusions, we remodel the IRST detection task as a one-dimensional signal anomaly detection task, and propose an efficient deep temporal probe network (DeepPro) that only performs calculations in the time dimension for IRST detection. We conducted extensive experiments to fully validate the effectiveness of our method. The experimental results are exciting, as our DeepPro outperforms existing state-of-the-art IRST detection methods on widely-used benchmarks with extremely high efficiency, and achieves a significant improvement on dim targets and in complex scenarios. We provide a new modeling domain, a new insight, a new method, and a new performance, which can promote the development of IRST detection. Codes are available at https://github.com/TinaLRJ/DeepPro.
@@ 1 Introduction
The objective of Infrared Small Target (IRST) detection [6, 31] is to identify and localize moving targets that are typically “small” and “dim”111“Small” means the target usually occupies a small portion of the overall image, and “dim” indicates that the target exhibits a weak signal intensity and low contrast relative to its background. from infrared images. It has been a longstanding, fundamental yet challenging problem in infrared remote sensing [74, 91], and has broad applicability in both civilian and military domains, including reconnaissance, precision guidance [68, 36], maritime surveillance [35, 27], early fire warnings [29, 110], vehicle detection [78, 66], and UAV monitoring [106, 30]. In recent years, IRST detection has attracted increasing attention and made measurable progress over the past few decades; however, most IRST detection methods have not been capable of simultaneously achieving outstanding performance in precision, robustness, and efficiency due to the following challenges [110, 10, 32, 90].
  • High-precision related challenges. On the one hand, the targets themselves are usually very small (e.g., typically occupy $\leq 9\times 9$ pixels), dim (e.g., with an SNR $\leq 3$) [36, 40, 19], and with insufficient appearance features, such as shape and texture information. Such characteristics increase the difficulty of developing effective feature representations. On the other hand, the background is usually nonsmooth and nonuniform, containing clutter and noise. Clutter refers to imaged real things, while noise refers to electronic noise and inaccuracies of the signal processor. They almost occupy the entire image, and some of them are strong and dynamic, similar to the targets as shown in Fig. 1(a) and (b). This complicates the effective suppression of false alarms.
  • High-robustness related challenges. Apart from the aforementioned challenges, the target characteristics evolve temporally under motion conditions. The intensity of the target varies, and the local background neighborhood of the target also changes dynamically. Furthermore, practical scenarios involve diverse complex backgrounds with distinct clutter properties and various targets with different grayscale distributions, as shown in Fig. 1(a) and (b). These complex factors make it difficult to detect targets continuously and stably.
@@
Therefore, in diverse complex scenarios, which modeling domain can better exhibit the targets? Which prior information can more effectively capture the essential differences between target and background? Exploring these questions is critical to achieve precise, robust, and efficient detection.
@@
Existing deep learning-based methods focus on extracting “more” information to improve the detection performance. For example, many single-frame methods [40, 95, 99] extensively explore the local saliency and global information in the spatial domain via the complex network structure. Besides, some multi-frame methods [18, 87, 70, 101] explore motion information in the short-term spatial-temporal domain. Although performance improvements are achieved in some scenarios, these methods suffer from severe performance degradation under complex conditions. This comes from the limitation that the target information is weak and that other discriminative information is insufficient in both domains, as illustrated in Fig. 1(c2) and (c3). In addition, extracting “more” information inevitably expands receptive fields, causing severe computational redundancy that fundamentally undermines real-time processing capabilities. Therefore, given our previous analysis, the focus of this work is to explore the “more essential” information from a more crucial profile and develop a highly effective and efficient method for IRST detection, especially under complex conditions.
@@
Figure 1: Visualizations of IRSTs in different scenarios and typical detection difficulties in different domains. (a, b) Distinct targets in different scenarios (Cases 1, 2, and 3): in appearance, targets and clutters show small inter-class differences, while different targets exhibit marked intra-class variations. (c1) Infrared image. (c2) Spatial domain: small dim targets are barely observable. (c3) Short-term spatial-temporal (ST) domain: targets are indistinguishable from interferences. (c4) Temporal profile: records dynamic statistical details of all signals at fixed spatial location, where small dim target signals are complete and more prominent. The temporal profile contains more essential information. (d, e) More contrasts of the three cases. For clarity, targets are marked with red annotations.
@@
In this paper, we study the inherent characteristics of IRST detection task in depth. We reveal that the temporal profile is crucial especially under complex conditions, which dramatically amplifies target saliency as shown in Fig. 1(c4) and (e). Through further theoretical analysis, we find that the temporal profile contains global temporal saliency information of target signals and correlation information between different signals, which are more essential for IRST detection. To validate the importance of the temporal profile in well-trained networks, we built the first prediction attribution [69, 86] tool that can analyze which input pixels impact the predictions of IRST detection networks. The attribution results show that the most influential pixels are concentrated in a small area centered on the target and continuously distributed along the time axis, which illustrates the importance of the temporal profile.
@@
Inspired by the above research, we remodel the IRST detection task as a one-dimensional signal anomaly detection task in the temporal profile, where targets manifest as statistical outliers in sequential signals. Then, we design a pixel-wise temporal probe mechanism (TPro) to extract essential information in the temporal profile. Based on TPro, we propose an effective and efficient deep temporal probe network (i.e., DeepPro) for IRST detection, which performs multiplications and additions only in the time dimension. We conducted numerous experiments on widely-used benchmarks. As shown in Fig. 2, our DeepPro achieves significantly superior performance as compared to existing state-of-the-art methods employing complex spatial-temporal computations. Meanwhile, our DeepPro exhibits higher robustness on dim targets and in complex scenarios with strong noise, achieving a 78.4% reduction in parameter count and over $2\times$ acceleration improvement compared to the lightest and fastest single-frame IRST (SIRST) detection network (i.e., Res-UNet [85]).
@@
  • Inspired by our research, we remodel the IRST detection task as a one-dimensional signal anomaly detection task. Then, we propose DeepPro to leverage the essential temporal profile information with multiply-add operations only in the time dimension.
  • Experimental results show that DeepPro achieves a significant performance improvement in IRST detection with extremely high efficiency, and has high robustness to dim targets and scenes with strong noise.
@@ 2.1 Single-frame infrared small target detection
Traditional SIRST detection methods mainly include spatial-filtering-based methods [62, 17, 3, 114], sparsity-and-low-rank-based methods [21, 15, 97, 111], and human-visual-system-based (HVS) methods [8, 75, 39, 89]. Recently, deep learning-based methods achieved better detection performance than traditional ones due to their strong feature learning abilities. Drawing inspiration from traditional methods, some networks are integrated with some prior knowledge for robust SIRST detection, such as the local spatial saliency differences between targets and backgrounds [14, 109, 88], the sparsity of targets and the low rankness of backgrounds [113, 82, 41]. Dai et al. [14] incorporated local contrast measure [80] into the network. Zhang et al. [109] converted the local contrast measure into a nonlocal quadrature difference measure in deep feature space. Zhou et al. [113] developed a deep unrolling network to exploit the low-rankness of backgrounds and the sparsity of targets. Liu et at. [41] combined low rankness and local smoothness prior of IRSTs to better model the background. Wu et al. [82] formulated the IRST detection task as sparse target extraction, low-rank background estimation, and image reconstruction in a relaxed robust principle component analysis model. For performance improvements, some methods combine IRST properties with the attention mechanism [72] to learn correlations among pixels and pay attention to key features. Li et al. [34] cascaded channel and spatial attention module to enhance multi-level features. Zhang et al. [104] proposed a Taylor finite difference-inspired edge block and a two-orientation attention aggregation block to highlight the shape of the targets. Zhang et al. [105] introduced a context mixer decoder based on spatial and frequency attention to modulate the low-level features for feature representation enhancement. Liu et al. [40] adopted the self-attention mechanism of Transformer to learn the non-local correlation of image features. Beside feature extraction, Liu et al. [41] proposed a scale and location sensitive (SLS) loss function to boost detection performance. Zhang et al. [98] integrated textual information with image modalities to enhance IRSTD performance. Zhang et al. [103] proposed a wavelet-based pruning method for efficient detection. Moreover, some researchers investigated the weakly-supervised detection [92, 33, 102], in which the pseudo-masks are generated based on the local saliency of the IRST. SIRST detection is applied to locate the salient targets, but is not suitable for targets with low SNR.
@@ 2.2 Multi-frame infrared small target detection
MIRST detection aims at exploiting both spatial and temporal information from input image tensors, and is expected to achieve better and more robust performance than SIRST detection. Existing MIRST detection methods can be categorized into filtering-based methods [55, 56, 22, 37, 71, 38, 51], sparsity-and-low-rank-based methods [107, 73, 45, 81, 35, 47, 43], and deep learning-based methods [67, 87, 36, 70, 100, 84, 79].
Filtering-based methods are the earlier developed methods, which mainly uses the temporal information among frames. Reed et al. [55, 56] proposed a series of 3D matched filtering methods to highlight the targets through spatial filtering and temporal energy accumulation. Gao et al. [22] proposed a time variance and patch contrast filter to improve the local contrast of targets. Li et al. [37] achieved MIRST detection by using the directional morphological filtering and spatial-temporal semantic information (e.g., the consistency of targets and the fluctuation of clutters). Mooney et al. [49] and Silverman et al. [60] early recognized the effectiveness of temporal profile filtering for detecting weak point IRSTs, making the “temporal profile” concept popular in IRST detection. Building on this foundation, some temporal profile-based filters [71, 38, 51] have been developed. These filters offer advantages in both detection performance and computational cost compared to other filtering-based methods.
@@ 3.1.1 Global temporal saliency
We analyzed the waveform characteristics of target signals in the temporal profile. Given that the velocity of small targets exhibits minimal variation over the short time they pass through a detection unit completely, we modeled the moving targets in uniform linear motions. In Fig. 3, we display some toy examples of the temporal profiles of the targets with different intensities, sizes, and speeds. The fluctuation of a target signal $G$ is regular with a single up-down process. As a target moves into and out of a detection unit, the target intensity in this unit first increases and then decreases. The amplitude of $G$ is related to the maximum intensity of the part of the target passing through the detection unit, and a brighter target usually has a higher peak value. Besides, the width of $G$ is related to the size and speed of the target. Bigger targets typically exhibit wider signals in the temporal profile, while targets with higher speeds under the same conditions produce narrower signals. Therefore, assuming that the fluctuations of other signals can be ignored, the target signals are saliency in the temporal profile.
Figure 4: Spatial and temporal-profile visualizations of targets in real scenes with noise and clutter. The red annotations mark the targets.
@@ 3.1.2 Correlation information
As the intensities of noise and clutter signals increase, the global temporal saliency becomes weak and damaged (e.g., Fig. 4(d)). Therefore, we deeply extract other essential distinctive characteristics between noise NN, clutter $B$, and target $G$ in the temporal profile.
@@
$\begin{array}[]{rl}R_{G,N}(\tau)&=\int_{-\infty}^{+\infty}f_{G+N}(t)\cdot f_{G+N}(-t+\tau)dt\\ &=f_{G+N}(\tau)*f_{G+N}(\tau).\end{array}$ (1)
@@
Then, to investigate the differences between target and clutter, we mixed target and different clutters (e.g., land clutter and moving cloud clutter), and analyzed the correlation function of the mixed signals (i.e., $f_{G+B}(t)$). At first, we got different clutters from the real video scenarios. After decentralizing and scaling, we added the fixed-intensity target signal to a series of clutter signals with different standard deviations $\sigma_{B}$ [^3] Note that the standard deviation of clutter in the time dimension is different from its fluctuation intensity in the space dimension. Usually, the temporal standard deviation of a clutter is smaller than its local spatial fluctuation since the background is slowly changing., as shown in Fig. 5(b1) and (c1). The typical clutters usually include low-frequency signals (e.g., slowly-changing cloud in c1) and high-frequency signals (e.g., camera blind elements and flickering background in b1). High-frequency signals are more likely to be falsely detected as targets due to their spatial and temporal saliency. When the clutter signal in the time dimension is not strong, there are some obviously distinctive characteristics between targets and clutters. Usually, the fluctuations of the clutter are dense, and they are correlated with each other on the amplitude in a period but uncorrelated on the occurrence moment. However, these characteristics are still inadequate for the detection when the clutter has a high standard deviation. Following the above analysis, we calculated the correlation function $R_{G,B}(\tau)$ of $f_{G+B}(t)$, i.e.,
$\begin{array}[]{rl}R_{G,B}(\tau)&=\int_{-\infty}^{+\infty}f_{G+B}(t)\cdot f_{G+B}(-t+\tau)dt\\ &=f_{G+B}(\tau)*f_{G+B}(\tau).\end{array}$ (2)